For the reaction A(aq)+B(aq)----- C(aq)+D(aq) the equilibrium constant is 24.6 at 25 c and 37.6 at 50 c What is the value of satndard Gibbs free energy change (in Kj) of this reaction at 75 c


1
Expert's answer
2021-06-10T03:34:13-0400

ΔG=VΔP−SΔT.

∆T= 50 - 25 = 25°C

= 37.6 - 24.6 = 13

= 0.75 *13 - 0.109*25

= 9.75 -2.725

= 7.025